Zander sold thrice as much face shield in the afternoon as in the morning. If he sold 340 pieces of face shield that day, how many face shields did he sell in the afternoon?
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39617)

What is confusing you about this exercise?

Answer by ikleyn(52781)

.

x in the morning

3x in the afternoon.


    x + 3x = 340

      4x   = 340

       x   = 340/4 = 85.


ANSWER.  85 in the morning  and 3*85 = 255 in the afternoon.

Solved.